# Reinforced MLLM: A Survey on RL-Based Reasoning in Multimodal Large Language Models

### v1

@@ 4.2.1. Outcome-supervised reward Mechanism
In addition to adopting the standard R1 training paradigm, Curr-ReFT (Deng et al., 2025b) further enhances reward alignment and output interpretability. Specifically, it incorporates a rejection sampling–based self-improvement mechanism that filters high-quality samples to reinforce foundational visual and linguistic competencies. Moreover, Curr-ReFT (Deng et al., 2025b) enforces output format constraints through structured reasoning templates, such as ¡thinking¿ and ¡answer¿ segments, ensuring strict alignment between intermediate reasoning processes and final answers. This output structuring improves semantic traceability, stabilizes reward feedback, and enhances the interpretability of reasoning stages, thereby extending the standardized R1 paradigm to more complex multimodal settings.
@@ 4.2.2. Process-Supervised Reward Mechanism
To mitigate sparse feedback and improve structural reasoning, recent studies adopt process-based rewards to enhance logical coherence. R1-VL (Zhang et al., 2025) serves as a representative example of this direction. It introduces Step-wise Group Relative Policy Optimization (StepGRPO), an online reinforcement learning framework that incorporates two structurally grounded reward components designed to operate over the reasoning process: StepRAR evaluates whether the model’s output includes key intermediate reasoning steps, while StepRVR assesses the logical coherence and completeness of the overall reasoning chain. These rewards are anchored at intermediate states and span the entire vision–language–reasoning trajectory, significantly improving the model’s ability to maintain logical consistency in complex tasks. Importantly, this structure-aware reward framework exhibits strong hierarchicality, as it captures not only final answers but also the internal logic layers and dependency structure within the reasoning process.
@@ 4.3.1. Curriculum Learning
Curriculum learning is a training strategy that emulates the human learning process by organizing tasks in an order of increasing difficulty, allowing models to first acquire basic skills before advancing to more complex reasoning abilities. In RL, especially for multimodal reasoning tasks, curriculum learning helps mitigate training instability and accelerates convergence by providing structured, progressively challenging learning stages. Kimi K1.5 employs curriculum sampling to progressively train the model from easy to difficult tasks, while integrating prioritized sampling to focus on the model’s weak areas. This approach effectively enhances training efficiency and overall performance during the curriculum learning process. Both strategies fully leverage prior information about problem difficulty and model performance, optimizing the learning process. Curr-ReFT introduces a curriculum-based reinforcement learning strategy that partitions training into three stages: binary classification, multiple-choice selection, and open-ended question answering. Each stage is guided by a task-specific reward function, enabling the model to gradually develop robust and scalable reasoning capabilities as task complexity increases.
@@ 4.3.2. Sample Efficiency
RL-based reasoning heavily relies on high-quality samples to guide effective policy optimization. Without sufficiently informative and reliable samples, the learning process may converge slowly or suboptimally, ultimately impairing the reasoning performance and generalization ability. Reason-RFT (Tan et al., 2025) first uses GPT-4o to filter out low-quality or incorrect samples and then reconstructs the filtered dataset to ensure its quality and suitability. Skywork R1V (Yi Peng, 2025) significantly enhances sample efficiency through Adaptive-Length Chain-of-Thought Distillation and a Hybrid Optimization Framework. The former dynamically adjusts the length of the reasoning chain to avoid excessive computation, while the latter iteratively optimizes the model by leveraging high-confidence samples and error samples, reducing the reliance on large-scale annotated data. Video-R1 (Feng et al., 2025)’s work on high-quality samples is primarily reflected in the construction of two datasets: Video-R1-COT-165k  [^1] https://huggingface.co/datasets/Video-R1/Video-R1-data and Video-R1-260k. These datasets contain carefully filtered image and video reasoning samples, ensuring the high quality of the data. MM-Eureka proposes a two-stage training strategy, with the second stage focusing on enhancing the model using domain-specific data. Additionally, the authors introduce the MMK12  [^2] https://huggingface.co/datasets/FanqingM/MMK12 multimodal mathematical reasoning dataset to highlight the importance of high-quality data. GFlowVLM (Kang et al., 2025) adopts generative flow networks to sample and learn from high-value reasoning trajectories, enabling efficient off-policy training and promoting diverse yet meaningful exploration. LMM-R1 (Peng et al., 2025) proposes a two-stage approach, FRE and MGT. In the first stage, the model’s reasoning ability is enhanced by training on purely textual data, and in the second stage, it generalizes to other modalities. This method effectively leverages high-quality textual data while alleviating the issue of scarcity in high-quality multimodal reasoning data. MetaSpatial (Pan and Liu, 2025) proposes a multi-turn refinement strategy based on GRPO, where grouped trajectories are evaluated to select higher-quality refinements that accelerate spatial reasoning acquisition.
@@ 4.3.3. Catastrophic forgetting
Catastrophic forgetting refers to the phenomenon where models, after fine-tuning on new tasks, lose previously acquired abilities. SFT often induces catastrophic forgetting by overwriting prior capabilities, whereas RL, by optimizing reward signals rather than explicit outputs, better preserves general reasoning skills (Liu et al., 2025c; Peng et al., 2025). In addition to the inherent advantages of RL, the mainstream solution in current RL-based methods to mitigate this issue is to incorporate a KL regularization term, which helps prevent excessive parameter drift (Pan et al., 2025; Liu et al., 2025e; Team et al., 2025). Curr-ReFT (Deng et al., 2025b) proposes Rejected Sample-based Self-improvement, a method that selectively learns from high-quality multimodal and textual examples to maintain the fundamental capabilities of MLLMs, thereby alleviating catastrophic forgetting. This approach ensures that while enhancing the model’s ability to perform new tasks, it retains its original knowledge and skills.
@@ 4.4. RLHF Training Framework
R1-V (Chen et al., 2025) MLLM GRPO
@@
Table 4. Overview of reinforcement learning from human feedback frameworks with supported models and RL algorithms.
In recent years, with the rapid development of reinforcement learning from human feedback (RLHF) and reward learning algorithms, a number of open-source libraries have emerged to provide implementation and experimentation tools, significantly lowering the research barrier and improving development efficiency. We summarize the currently popular open-source RLHF training frameworks in Table 4 and present the RL algorithms they currently support. DeepSpeed-Chat (Rasley et al., 2020) focuses on training ChatGPT-like models, enabling large-scale model training through RLHF in a fast and cost-effective manner while reducing training time and expenses. OpenRLHF (Hu et al., 2024), built on Ray (Moritz et al., 2018), DeepSpeed, and HuggingFace Transformers (Wolf et al., 2020), supports distributed training and various algorithms (e.g., PPO (Schulman et al., 2017), REINFORCE++ (Hu, 2025)), optimized for performance and memory usage in ultra-large models (70B+ parameters), and remains compatible with the HuggingFace ecosystem for multi-domain tasks. LLaMA-Factory (Zheng et al., 2024) supports fine-tuning of over a hundred large models, covering various training methods and precision optimization strategies, and enhances the development experience with a visual dashboard and efficient inference interfaces, making it suitable for scenarios requiring broad task coverage and high performance. TRL (von Werra et al., 2020) based on Transformers, focuses on post-training techniques such as SFT, PPO (Schulman et al., 2017), and DPO (Rafailov et al., 2023), offering efficient distributed training support and versatile trainer interfaces for large-scale model optimization across different hardware environments. Colossal-AI (Li et al., 2023) simplifies the development of distributed deep learning models by providing multiple parallel strategies and user-friendly tools, significantly reducing resource costs and supporting practical applications such as video generation and large language model fine-tuning. VeRL (Sheng et al., 2024) is a flexible and efficient reinforcement learning library that supports multiple algorithms and integrates with mainstream frameworks, with a modular design and advanced performance optimization techniques that make it well-suited for training multimodal and vision-language models. Open-R1 (Face, 2025) reproduces the complete pipeline of DeepSeek-R1, offering scripts for training, evaluation, and synthetic data generation, along with step-by-step guides to help users get started quickly. R1-V (Chen et al., 2025) focuses on reinforcement learning and long-CoT-based reasoning in vision-language models, driving the development of general visual-language intelligence through algorithmic enhancements and diverse task designs. EasyR1 (Yaowei Zheng, 2025), a clean fork of OpenR1, is a multimodal reinforcement learning framework featuring a hybrid engine design, support for checkpoint resumption, and seamless integration with tools like Wandb. As RLHF research involves complex human interactions and multi-component systems, implementing new ideas and conducting experiments can be challenging; however, these frameworks offer diverse options with unique strengths, accelerating innovation and progress in the field.
@@ 5.1. Embodied AI
Embodied AI focuses on enabling MLLMs to perceive, reason, and act in physical or simulated environments. RL-based multimodal reasoning methods have become pivotal in improving models’ ability to interpret multimodal inputs and to generate action-oriented reasoning strategies. Models such as MetaSpatial (Pan and Liu, 2025) demonstrate the capability of integrating physics-based rewards and perceptual feedback to support three-dimensional scene understanding and object manipulation. Similarly, Video-R1 (Feng et al., 2025) extends reasoning into the temporal domain, enabling video-based agents to infer causal relationships between sequential frames, a critical aspect for embodied perception in dynamic environments. The reinforcement learning framework facilitates iterative learning of spatial, temporal, and causal relations, allowing embodied MLLMs to achieve stronger generalization and physical realism.
@@ 5.2. MLLM Agentic System
Recent RL-based methods have made substantial progress toward building agentic MLLMs. R1-Onevision (Yang et al., 2025) pioneers a two-stage approach, first constructing CoT datasets enriched with structured image descriptions to ground the model’s reasoning in perceptual inputs, and then applying RL to refine autonomous reasoning capabilities. This design ensures that agents can actively retrieve and integrate visual information throughout their decision process, a core requirement for embodied agentic behavior. Skywork R1V (Yi Peng, 2025) further advances this direction by proposing a lightweight agentic training paradigm. Instead of fine-tuning the entire model, it iteratively updates a vision projection module atop a reasoning-capable language backbone, significantly reducing hardware costs while enabling scalable vision-language interaction for agentic tasks. This strategy reflects a trend toward modular agentic systems that efficiently align perception and reasoning components. In interactive scenarios such as GUI-based task execution, UI-R1 (Lu et al., 2025) demonstrates the importance of multimodal action reasoning. By jointly optimizing rewards for action type prediction, argument selection, and output formatting, the model effectively learns to perform sequences of operations, aligning its outputs with human-intended goals. This illustrates the shift from static prediction to dynamic interaction—an essential hallmark of agentic MLLMs.
@@ 5.3. Domain-Specific Applications of RL-Based Multimodal Reasoning
Beyond general embodied and agentic capabilities, reinforcement learning–based multimodal reasoning has been increasingly applied to a range of specialized real-world domains that demand high levels of perception, reasoning, and decision-making under complex conditions. By leveraging structured rewards and sequential learning frameworks, these methods enable MLLMs to move beyond static understanding toward dynamic, context-sensitive behavior. Domains such as autonomous driving, medical and healthcare, and human-centered interaction serve as critical testbeds for evaluating and advancing the robustness, generalization, and interpretability of RL-empowered multimodal systems.
5.3.1. Autonomous Driving
Autonomous driving represents a highly challenging domain for reasoning, requiring the integration of visual perception, temporal understanding, spatial reasoning, and real-time decision-making. Although explicit large-scale RL-based MLLM applications to driving are still emerging, foundational work such as Video-R1 (Feng et al., 2025) and MetaSpatial (Pan and Liu, 2025) has laid critical groundwork. Video-R1 (Feng et al., 2025) introduces a temporal reasoning mechanism that rewards models for maintaining logical consistency across video frames, simulating the sequential perception process needed for understanding dynamic traffic scenes. This capacity to model temporal causality is crucial for driving agents tasked with interpreting pedestrian movements, vehicle trajectories, and traffic signal changes. Similarly, MetaSpatial (Pan and Liu, 2025) demonstrates the feasibility of spatial reasoning through structured physics-based rewards, ensuring that models can reason about object interactions, collision avoidance, and environment constraints—skills that are directly transferable to autonomous vehicle navigation. Reinforcement learning plays a key role in these systems by enabling continuous adaptation to novel environments, unexpected events, and evolving road scenarios, pushing MLLMs toward real-world, agentic driving intelligence.
@@ 5.3.2. Medical and Healthcare
The medical and healthcare domains present unique demands for high-stakes reasoning, interpretability, and generalization. RL-based multimodal reasoning methods have shown promising advancements in this area, particularly in medical visual question answering and clinical decision support. MedVLM-R1 (Pan et al., 2025) exemplifies how the adaptation of verifiable reward mechanisms can guide models to generate explicit reasoning paths for medical multiple-choice questions, significantly improving both in-domain accuracy and out-of-distribution robustness. By rewarding models not only for correct final answers but also for structured reasoning traces, RL training enhances model transparency—an essential attribute for healthcare applications where decision rationales must be verifiable by medical professionals. Moreover, the reinforcement of structured outputs helps reduce hallucinations and ensures that medical MLLMs remain faithful to diagnostic evidence presented in multimodal inputs such as X-rays, CT scans, and clinical notes. Looking forward, RL-based methods offer a pathway toward more reliable, explainable, and adaptable AI assistants in healthcare, supporting clinicians across a wide range of diagnostic and treatment planning tasks.
@@ 5.3.3. Social and Human
Understanding human behavior, emotions, and social interactions is another frontier where RL-based multimodal reasoning plays an increasingly important role. Models such as R1-Omni (Zhao et al., 2025) explore emotion recognition by integrating audio, video, and textual modalities, leveraging reinforcement learning to fine-tune reasoning strategies that capture subtle human affective cues. By jointly optimizing for emotional classification and reasoning structure, RL enables models to move beyond surface-level perception, fostering deeper socio-emotional understanding. In parallel, Qwen2-Audio-Instruct-RL (Li et al., 2025a) highlights the potential of auditory reasoning, training MLLMs to interpret acoustic signals such as tone, speech dynamics, and environmental sounds within a broader reasoning framework. These capabilities are critical for developing socially aware dialogue systems, assistive agents for elderly care, and empathetic AI companions. Reinforcement learning enhances the robustness and adaptability of such systems by enabling dynamic interaction modeling, error recovery during conversations, and personalized adaptation to individual users’ communicative styles. As MLLMs evolve into socially intelligent agents, RL-based multimodal reasoning will be pivotal for achieving nuanced, context-sensitive human-AI interaction.
@@ 6. Datasets and Benchmark
Benchmark Reasoning over Structured Knowledge Domains Contextual and Perceptual Reasoning in Real-World Scenarios Generalization and Multi-Task Reasoning Benchmarks Mathematical Chart Scientific Code Generation Algorithmic Problem Social and Cultural Knowledge WE-MATH (Qiao et al., 2024); Mathvista (Lu et al., 2023); MathVerse (Zhang et al., 2024b); MATH-Vision (Wang et al., 2024b); MathScape (Zhou et al., 2024); CMM-Math (Liu et al., 2024c); OlympiadBench (He et al., 2024); MV-MATH (Wang et al., 2025a); GeoSense (Xu et al., 2025c) ChartBench (Xu et al., 2023); MultiChartQA (Zhu et al., 2024); Charxiv (Wang et al., 2024c) M4U (Wang et al., 2024d); MMMU (Yue et al., 2024a); MMMU-Pro (Yue et al., 2024b); ScienceQA (Lu et al., 2022); TheoremQA (Chen et al., 2023a); EMMA (Hao et al., 2025); GAOKAO-MM (Zong and Qiu, 2024); CMMMU (Ge et al., 2024); MM-Star (Chen et al., 2024b) ChartMimic (Shi et al., 2024); Plot2Code (Wu et al., 2024b); HumanEval-V (Zhang et al., 2024c) NPHardEval4V (Fan et al., 2024) II-Bench (Liu et al., 2024a); CII-Bench (Zhang et al., 2024a); PunchBench (Ouyang et al., 2024); GeoComp (Song et al., 2025) Action Prediction Spatial Reasoning InfiGUIAgent-Data (Liu et al., 2025b); Mind2Web (Deng et al., 2023); SeeClick (Cheng et al., 2024); Driving-Hazard-Prediction-and-Reasoning (Charoenpitaks et al., 2024); iVISPAR (Mayer et al., 2025); PulseCheck457 (Wang et al., 2025b); VisuLogic (Xu et al., 2025b) Multi-Image Based Inductive Reasoning Other Composite Benchmarks MM-IQ (Cai et al., 2025); LogicVista (Xiao et al., 2024); PuzzleTest (Toh et al., 2025) VRC-Bench (Thawakar et al., 2025); M3CoT (Chen et al., 2024d); MLLM-COMPBench (Kil et al., 2024); MME-CoT (Jiang et al., 2025); ZeroBench (Roberts et al., [n. d.])
@@
To comprehensively evaluate the reasoning capabilities of MLLMs, researchers have developed a variety of benchmark datasets spanning different reasoning paradigms and modalities. In this section, we categorize these datasets into three high-level domains: Reasoning over Structured Knowledge Domains, Contextual and Perceptual Reasoning in Real-World Scenarios, and Generalization and Multi-Task Reasoning Benchmarks, with each domain encompassing several sub-domains that reflect distinct types of knowledge, reasoning structures, or contextual challenges. An overview of these representative benchmarks is illustrated in Figure 3. Among them, a significant portion of current evaluations is concentrated in the first domain, particularly in mathematical, scientific, and chart-based reasoning tasks. We summarize the performance of current open-source and closed-source reasoning models on mainstream reasoning benchmarks in Table 5. Among the open-source models, Mulberry (Yao et al., 2024), LLaVA-CoT (Xu et al., 2024), Insight-V (Dong et al., 2024), and LLaVA-Reasoner (Zhang et al., 2024d) are non-RL-based models, while the remaining models are RL-based. However, this focus also reveals a fundamental limitation: the dominance of well-structured, mathematically grounded benchmarks leaves open the question of whether these models can generalize to more complex, open-ended, and perceptually rich scenarios. Tasks that require interaction with dynamic environments, spatial abstraction, multimodal alignment, or socio-cultural reasoning are underrepresented, despite being critical for real-world applications. To address this limitation, we further include a range of benchmarks under the latter two categories, covering diverse scenarios such as user interface action prediction, 6D spatial reasoning, visual-to-code generation, and culturally grounded multimodal understanding. These datasets provide a more comprehensive lens for evaluating MLLMs’ reasoning abilities beyond traditional symbolic domains. In the following, we present a detailed analysis of representative benchmarks within each sub-domain.
Model MathVista MathVision MathVerse MMMU-Val MM-Star MME ChartQA DynaMath HallBench Olympiad
OpenAI o1 71.00 - - 77.30 - - - - - -
GPT-4o 0513 63.80 30.40 39.40 69.10 63.90 2329.00 85.70 63.70 55.00 -
Claude-3.5-snooet-1022 65.30 35.60 - 66.40 62.20 1920.00 90.80 64.80 55.00 -
QVQ-72B-Preview 71.40 35.90 - 70.30 - - - - - -
Mulberry-7B 63.1 - - 55.0 61.3 2396 83.9 45.1 54.1 -

### v2

@@ 4.2.1. Outcome-supervised reward Mechanism
Model Reward
Embodied-R (Zhao et al., 2025a) $r_{i}=\omega_{1}\cdot r^{\text{format}}_{i}+\omega_{2}\cdot r^{\text{accuracy}}_{i}+\omega_{3}\cdot r^{\text{logic}}_{i}$
NoisyRollout (Liu et al., 2025d) $R_{i}=\begin{cases}1,&\mathrm{if~{}}o_{i}\mathrm{~{}is~{}correct}\\ 0,&\mathrm{otherwise}\end{cases}$, $\bar{R}=\frac{1}{n_{1}+n_{2}}\sum_{i=1}^{n_{1}+n_{2}}R_{i}$
Table 5. Reward function of RL-Based MLLM reasoning methods for Curriculum-Based Reward Strategy.
In addition to adopting the standard R1 training paradigm, Curr-ReFT (Deng et al., 2025b) employs a three-stage curriculum consisting of binary classification, multiple-choice selection, and open-ended question answering, with each stage guided by task-specific rewards to progressively build reasoning ability. Embodied-R (Zhao et al., 2025a) also adopts a staged training approach, introducing different weights for $\omega$ across training phases. Additionally, it incorporates a logical consistency reward to align the reasoning process with the final answer in spatial tasks. NoisyRollout (Liu et al., 2025d) enhances exploration by mixing clean and noisy visual inputs, using a noise annealing schedule to guide training from early diversity to stable convergence.
@@ 4.2.2. Process-Supervised Reward Mechanism
To mitigate sparse feedback and improve structural reasoning, recent studies adopt process-based rewards to enhance logical coherence. R1-VL (Zhang et al., 2025a) serves as a representative example of this direction. It introduces Step-wise Group Relative Policy Optimization (StepGRPO), an online reinforcement learning framework that incorporates two structurally grounded reward components designed to operate over the reasoning process: StepRAR evaluates whether the model’s output includes key intermediate reasoning steps, while StepRVR assesses the logical coherence and completeness of the overall reasoning chain. These rewards are anchored at intermediate states and span the entire vision–language–reasoning trajectory, significantly improving the model’s ability to maintain logical consistency in complex tasks. Similarly, ChestX-Reasoner (Fan et al., 2025) introduces a process reward based on clinical reports, evaluating the factual accuracy of intermediate reasoning steps to enhance diagnostic coherence and reasoning quality.
@@ 4.3.1. Curriculum Learning
Curriculum learning is a training strategy that emulates the human learning process by organizing tasks in an order of increasing difficulty, allowing models to first acquire basic skills before advancing to more complex reasoning abilities. In RL, especially for multimodal reasoning tasks, curriculum learning helps mitigate training instability and accelerates convergence by providing structured, progressively challenging learning stages. Kimi K1.5 (Team et al., 2025a) adopts curriculum and prioritized sampling to expose the model to increasingly difficult examples while focusing on weak areas, enhancing overall training efficiency. Curr-ReFT (Deng et al., 2025b) introduces difficulty-aware reward shaping to stage the learning process, facilitating stable reward-based training. Beyond data and reward curricula, Embodied-R (Zhao et al., 2025a) implements a three-stage RL schedule that gradually shifts reward weights from format to accuracy and logical consistency, guiding the model toward coherent embodied reasoning. NoisyRollout (Liu et al., 2025d) implicitly applies curriculum via a noise annealing schedule—starting with distorted visual inputs and gradually reducing noise—promoting early exploration and stable convergence without added cost.
@@ 4.3.2. Sample Efficiency
RL-based reasoning heavily relies on high-quality samples to guide effective policy optimization. Insufficient or unreliable samples can lead to slow convergence and suboptimal performance, ultimately impairing reasoning and generalization. Skywork R1V (Yi Peng, 2025) improves sample efficiency via Adaptive-Length CoT Distillation, which adjusts reasoning chain length to reduce computation, and a Hybrid Optimization Framework that iteratively refines the model using both high-confidence and error samples, reducing dependence on large-scale annotations. MM-Eureka (Meng et al., 2025) proposes a two-stage training strategy, with the second stage focusing on enhancing the model using domain-specific data. Additionally, the authors introduce the MMK12  [^1] https://huggingface.co/datasets/FanqingM/MMK12 multimodal mathematical reasoning dataset to highlight the importance of high-quality data. GFlowVLM (Kang et al., 2025) adopts generative flow networks to sample and learn from high-value reasoning trajectories, enabling efficient off-policy training and promoting diverse yet meaningful exploration. LMM-R1 (Peng et al., 2025) employs a two-stage method, FRE followed by MGT, where the model is first trained on textual data to enhance reasoning and then generalized to other modalities, thereby mitigating the scarcity of high-quality multimodal data. MetaSpatial (Pan and Liu, 2025) introduces a GRPO-based multi-turn refinement strategy, selecting higher-quality grouped trajectories to accelerate spatial reasoning. NoisyRollout (Liu et al., 2025d) demonstrates strong generalization with only 2.1K samples by mixing clean and noisy rollouts, enhancing learning through visual diversity at no extra cost.
@@ 4.3.3. Catastrophic forgetting
Catastrophic forgetting refers to the phenomenon where models, after fine-tuning on new tasks, lose previously acquired abilities. SFT often induces catastrophic forgetting by overwriting prior capabilities, whereas RL, by optimizing reward signals rather than explicit outputs, better preserves general reasoning skills (Liu et al., 2025e; Peng et al., 2025). In addition to the inherent advantages of RL, the mainstream solution in current RL-based methods to mitigate this issue is to incorporate a KL regularization term, which helps prevent excessive parameter drift (Pan et al., 2025; Liu et al., 2025g; Team et al., 2025a). Curr-ReFT (Deng et al., 2025b) proposes Rejected Sample-based Self-improvement, a method that selectively learns from high-quality multimodal and textual examples to maintain the fundamental capabilities of MLLMs. This approach ensures that while enhancing the model’s ability to perform new tasks, it retains its original knowledge and skills.
@@ 4.4. RLHF Training Framework
R1-V (Chen et al., 2025a) MLLM GRPO
@@
Table 6. Overview of RLHF frameworks with supported models and RL algorithms.
In recent years, the rise of RLHF and reward learning has led to the emergence of various open-source libraries that lower research barriers and enhance development efficiency. Table 6 summarizes major RLHF frameworks and their supported algorithms. DeepSpeed-Chat (Rasley et al., 2020) enables efficient, cost-effective ChatGPT-style model training. OpenRLHF (Hu et al., 2024) supports distributed training for large models and integrates with HuggingFace. LLaMA-Factory (Zheng et al., 2024) offers broad model and training method support with a visual interface. TRL (von Werra et al., 2020) focuses on post-training techniques like PPO and DPO with distributed support. Colossal-AI (Li et al., 2023) simplifies large-scale training with parallel strategies. VeRL (Sheng et al., 2024) targets multimodal and vision-language tasks with modular design. Open-R1 (Face, 2025) and its fork EasyR1 (Yaowei Zheng, 2025) replicate and extend the DeepSeek-R1 pipeline. R1-V (Chen et al., 2025a) emphasizes long-chain reasoning in vision-language models. Together, these frameworks offer diverse capabilities that accelerate innovation in RLHF research.
@@ 5.1. Embodied AI
Embodied AI focuses on enabling MLLMs to perceive, reason, and act in physical or simulated environments. RL-based multimodal reasoning methods have become pivotal in improving models’ ability to interpret multimodal inputs and to generate action-oriented reasoning strategies. Models such as MetaSpatial (Pan and Liu, 2025) demonstrate the capability of integrating physics-based rewards and perceptual feedback to support three-dimensional scene understanding and object manipulation. Similarly, Video-R1 (Feng et al., 2025) extends reasoning into the temporal domain, enabling video-based agents to infer causal relationships between sequential frames, a critical aspect for embodied perception in dynamic environments. Embodied-R (Zhao et al., 2025a) proposes a collaborative RL framework that activates embodied spatial reasoning by decoupling perception and reasoning, achieving superior performance and generalization with a logical consistency reward. The reinforcement learning framework facilitates iterative learning of spatial, temporal, and causal relations, allowing embodied MLLMs to achieve stronger generalization and physical realism. Autonomous driving can also be viewed as a representative embodied scenario, where RL-enhanced MLLMs hold promise for future advances in spatial-temporal reasoning and decision-making.
@@ 5.2. MLLM Agentic System
In interactive scenarios such as GUI-based task execution, UI-R1 (Lu et al., 2025) demonstrates the importance of multimodal action reasoning. By jointly optimizing rewards for action type prediction, argument selection, and output formatting, the model effectively learns to perform sequences of operations, aligning its outputs with human-intended goals. GUI-R1 (Xia and Luo, 2025) enhances the GUI capabilities of MLLM in advanced real-world task scenarios by unifying action types, input text, and click coordinates into a standardized action space framework. InfiGUI-R1 (Liu et al., 2025c) advances GUI agents from reactive execution to deliberative reasoning via a two-stage RL framework, enhancing planning and error recovery through sub-goal guidance and reflective correction.
@@ 5.3. Domain-Specific Applications of RL-Based Multimodal Reasoning
Beyond general embodied and agentic capabilities, RL–based multimodal reasoning has been increasingly applied to a range of specialized real-world domains that demand high levels of perception and decision-making under complex conditions. By leveraging structured rewards and sequential learning frameworks, these methods enable MLLMs to move beyond static understanding toward dynamic, context-sensitive behavior. Domains such as medical and healthcare, as well as human-centered interaction, serve as critical testbeds for evaluating and advancing the robustness, generalization, and interpretability of RL-empowered multimodal systems.
@@ 5.3.1. Medical and Healthcare
The medical and healthcare domains present unique demands for high-stakes reasoning, interpretability, and generalization. RL-based multimodal reasoning methods have shown promising progress in tasks such as medical visual question answering and clinical decision support. MedVLM-R1 (Pan et al., 2025) improves answer verifiability by rewarding structured reasoning traces for multiple-choice tasks. ChestX-Reasoner (Fan et al., 2025) applies process-supervised reinforcement learning to align model reasoning with clinical workflows, using supervision signals extracted from radiology reports. This approach enhances factuality, completeness, and diagnostic relevance of generated reasoning chains, while also improving outcome accuracy across diverse tasks such as disease classification, anomaly detection, and temporal comparison.
@@ 5.3.2. Social and Human
Understanding human behavior, emotions, and social interactions is an emerging frontier for RL-based multimodal reasoning. R1-Omni (Zhao et al., 2025b) integrates audio, video, and text to enhance emotional recognition via RL, enabling structured socio-emotional reasoning. Similarly, R1-AQA (Li et al., 2025b) trains MLLMs to interpret acoustic signals for auditory reasoning. These capabilities support empathetic AI agents and socially adaptive systems, where RL facilitates robustness, error recovery, and personalized interaction.
@@ 6. Datasets and Benchmark
Benchmark Reasoning over Structured Knowledge Domains Contextual and Perceptual Reasoning in Real-World Scenarios Generalization and Multi-Task Reasoning Benchmarks Mathematical Chart Scientific Code Generation Algorithmic Problem Social and Cultural Knowledge WE-MATH (Qiao et al., 2024); Mathvista (Lu et al., 2023); MathVerse (Zhang et al., 2024b); MATH-Vision (Wang et al., 2024b); MathScape (Zhou et al., 2024); CMM-Math (Liu et al., 2024c); OlympiadBench (He et al., 2024); MV-MATH (Wang et al., 2025a) ChartBench (Xu et al., 2023); MultiChartQA (Zhu et al., 2024); Charxiv (Wang et al., 2024c) M4U (Wang et al., 2024d); MMMU (Yue et al., 2024a); MMMU-Pro (Yue et al., 2024b); ScienceQA (Lu et al., 2022); TheoremQA (Chen et al., 2023); EMMA (Hao et al., 2025); GAOKAO-MM (Zong and Qiu, 2024); CMMMU (Ge et al., 2024); MM-Star (Chen et al., 2024a); MDK12-Bench (Zhou et al., 2025b) ChartMimic (Shi et al., 2024); Plot2Code (Wu et al., 2024); HumanEval-V (Zhang et al., 2024c) NPHardEval4V (Fan et al., 2024) II-Bench (Liu et al., 2024a); CII-Bench (Zhang et al., 2024a); PunchBench (Ouyang et al., 2024); GeoComp (Song et al., 2025) Action Prediction Spatial Reasoning InfiGUIAgent-Data (Liu et al., 2025b); Mind2Web (Deng et al., 2023); SeeClick (Cheng et al., 2024); Driving-Hazard-Prediction-and-Reasoning (Charoenpitaks et al., 2024); iVISPAR (Mayer et al., 2025); PulseCheck457 (Wang et al., 2025b); VisuLogic (Xu et al., 2025b); GeoSense (Xu et al., 2025c) Multi-Image Based Inductive Reasoning Other Composite Benchmarks MM-IQ (Cai et al., 2025); LogicVista (Xiao et al., 2024); PuzzleTest (Toh et al., 2025) VRC-Bench (Thawakar et al., 2025); M3CoT (Chen et al., 2024c); MLLM-COMPBench (Kil et al., 2024); MME-CoT (Jiang et al., 2025); ZeroBench (Roberts et al., [n. d.]); VCR-Bench (Qi et al., 2025)
@@
To comprehensively evaluate the reasoning capabilities of MLLMs, researchers have developed a variety of benchmark datasets spanning different reasoning paradigms and modalities. In this section, we categorize these datasets into three high-level domains: Reasoning over Structured Knowledge Domains, Contextual and Perceptual Reasoning in Real-World Scenarios, and Generalization and Multi-Task Reasoning Benchmarks, with each domain encompassing several sub-domains that reflect distinct types of knowledge, reasoning structures, or contextual challenges. An overview of these representative benchmarks is illustrated in Figure 3. We summarize the performance of current open-source and closed-source reasoning models on mainstream reasoning benchmarks in Table 7. Among these, Mulberry (Yao et al., 2024), LLaVA-CoT (Xu et al., 2024), Insight-V (Dong et al., 2024), and LLaVA-Reasoner (Zhang et al., 2024d) are mainstream reasoning models that do not rely on RL, while the remaining models are RL-based. From Table 7, it is evident that a significant portion of current evaluations is concentrated in the first domain, particularly in mathematical, scientific, and chart-based reasoning tasks. This focus reveals a fundamental limitation: the dominance of well-structured, mathematically grounded benchmarks leaves open the question of whether these models can generalize to more complex, open-ended, and perceptually rich scenarios. Tasks that require interaction with dynamic environments, spatial abstraction, multimodal alignment, or socio-cultural reasoning are underrepresented, despite being critical for real-world applications. To address this limitation, we further include a range of benchmarks under the latter two categories, covering diverse scenarios such as user interface action prediction, 6D spatial reasoning, visual-to-code generation, and culturally grounded multimodal understanding. These datasets provide a more comprehensive lens for evaluating MLLMs’ reasoning abilities beyond traditional symbolic domains. In the following, we present a detailed analysis of representative benchmarks within each sub-domain.
Model MathVista MathVision MathVerse MMMU-Val MM-Star MME ChartQA DynaMath HallBench MMVet
OpenAI o1 71.0 - - 77.3 - - - - - -
GPT-4o 0513 63.8 30.4 39.4 69.1 63.9 2329.0 85.7 63.7 55.0 -
Claude-3.5-snooet-1022 65.3 35.6 - 66.4 62.2 1920.0 90.8 64.8 55.0 -
QVQ-72B-Preview 71.4 35.9 - 70.3 - - - - - -
Mulberry-7B 63.1 - - 55.0 61.3 2396.0 83.9 45.1 54.1 -